A collection of 15 coins is worth 1.00. If the coins are dimes and nickels, how many of each coin are there?
This is a word problem from Math 100.
Can you please give me the the equation for solving this problem?

Found 2 solutions by anne, tutorcecilia:
Answer by anne(47)

hi. I did a similar question to this too and solved it by this equations
n+d=1.00
0.1d+0.05n=1.00
and so on

Answer by tutorcecilia(2152)

Let x = nickles
Let y = dimes
.
.05x + .10y = 1.00
x+y=15
y=15-x
So,
.05x + .10(15-x)=1.00
.05x + 1.5 - .10x = 1.00
-.05x + 1.5 = 1.00
-.05x = 1.00 - 1.5
-.05x = -.50
-.05x/-.05 = -.50/-.05
x=10
x+y=15
10+y=15
y=15-10
y=5
.
Checking:
0.05(10) + .10(5)=1.00
.50 +.50=1.00
1.00=1.00 (checks out)